# The Learning Approach to Games

### v2

@@ Abstract
This work introduces a unified framework for a more detailed exploration of games. In existing literature, strategies of players are typically assigned scalar values, and the concept of Nash equilibrium is used to identify compatible strategies. However, this approach lacks the internal structure of a player, thereby failing to accurately model observed behaviors.
To address this limitation, we propose an abstract definition of a player. This allows for a more nuanced understanding of players and brings the focus to the challenge of learning that players face. Unlike Markov decision processes, which formalize control problems but not agent design, our framework subsumes standard reinforcement learning structures. It thus offers a language that enables a deeper connection between games and learning. To illustrate the need for such generality, we study a simple two-player game and show that even in the most basic settings, a sophisticated player may adopt dynamic strategies that cannot be captured by simpler designs or compatibility analysis alone.
In the discrete setting, we consider a player whose structure incorporates standard estimates from the literature. We explore connections to correlated equilibrium and highlight that dynamic programming naturally applies to all estimates. In the mean-field setting, we exploit symmetry to construct explicit examples of equilibria. Finally, we examine connections to reinforcement learning and bandit problems, demonstrating the broad applicability of the framework.
@@ 1 Introduction
Mathematically, we need to have fundamental concepts that are general enough to cover everything we aim to model. In the realm of games, we point out that what defines a player is a sequence of observations, a sequence of potential actions, and a collection of learning algorithms, all of which might be remarkably complicated. Given that the universe, or environment, is characterized by some abstract probability space $(\Omega^{u},{\cal F}^{u},\mathbb{P}^{u})$, and denoting ${\cal P}(E)$ as the set of probability distributions on the set $E$, we introduce the definition of a player as arbitrary as possible;
Definition 1
Let ${\cal E}$ be a space of observables, and $\mathscr{E}$ be the set of finite sequences of ${\cal E}$. Also, let $\mathbb{A}$ be a space of actions, and $\mathscr{A}$ be the set of finite sequences of $\mathbb{A}$. Moreover, let $\mathscr{M}_{1},\dots,\mathscr{M}_{k}$ be some arbitrary set of functions with domain ${\cal D}$, called spaces of estimations.
We call $({\cal O},\mathfrak{L}_{1},\cdots,\mathfrak{L}_{k},{\Upsilon})$ a player in the environment $(\Omega^{u},{\cal F}^{u},\mathbb{P}^{u})$ with observations ${\cal O}$, learning algorithms $\mathfrak{L}_{1},\dots,\mathfrak{L}_{k}$ and with behavior ${\Upsilon}$, if
${\cal O}:\Omega^{u}\times\mathbb{N}\to\mathscr{E}\ \ \text{satisfying}\ \ {}^{n}{\cal O}\ \ \text{is a subsequence of}\ \ {}^{n+1}{\cal O}\ \text{where}\ {}^{n}{\cal O}:={\cal O}(\omega^{u},n)$ (1.1)
for all $\ n\in\mathbb{N},\ w^{u}\in\Omega^{u}$;
$\mathfrak{L}_{\ell}:\mathscr{E}\to\mathscr{M}_{\ell},\quad\forall\ell\in\{1,\dots,k\}$
and
${\Upsilon}:\mathscr{M}_{1}\times\cdots\times\mathscr{M}_{k}\to({\cal D}\to{\cal P}(\mathscr{A}))$
Furthermore, we call
${}^{n}{\Upsilon}:\Omega^{u}\times\mathbb{N}\to({\cal D}\to{\cal P}(\mathscr{A})),\quad{}^{n}{\Upsilon}:={\Upsilon}(\mathfrak{L}_{1}({}^{n}{\cal O}),\dots,\mathfrak{L}_{k}({}^{n}{\cal O}))$
is the planned behavior of the player at age $n$.
Note that a common domain ${\cal D}$ is not restrictive and might have subdomains that each estimation depends on. As an example, it is standard to have a model with time and state space included in ${\cal D}$. Also, typically ${\cal E}\subset\mathbb{A}\times{\cal D}$. When there are more than one player, we keep track of them by the index $i\in\mathbb{N}_{0}:=\{1,2,\dots\}$, and set $\vec{\varphi}:=\prod_{i\in\mathbb{N}_{0}}\varphi^{i}$ for any notation in the Definition 1. In this case, typically ${\cal E}^{i}\subset\vec{\mathbb{A}}\times\vec{\cal D}$ and players are observing each other. As an example, there might be a common observed state space, or players might observe each other’s actions.
Now, equip the space $({\cal D}\to{\cal P}(\mathscr{A}))$ with a metric $d$ and define the concept of convergence as follows;
Definition 2
Given a player $({\cal O},\mathfrak{L}_{1},\dots,\mathfrak{L}_{k},{\Upsilon})$, we say ${\Upsilon}^{*}\in({\cal D}\to{\cal P}(\mathscr{A}))$ is a $(r,\delta)$-recurrent behavior if
$\mathbb{P}^{u}\Big{(}\operatorname*{liminf}_{n\to\infty}d({\Upsilon}^{*},{}^{n}{\Upsilon})>r\Big{)}\leq\delta$
Also, we say ${\Upsilon}^{*}$ is almost surely a recurrent behaviour of the player if $r=\delta=0$.
We can generalize the definition to more than one player in a straightforward manner. In this case, as they interact with each other, a collectively recurrent behavior forms the basis for an equilibrium.
Next, let us further assume that the spaces of estimations $\mathscr{M}_{1},\dots,\mathscr{M}_{k}$ are also equipped with metrics $d_{1},\dots,d_{k}$. Then, if ${\Upsilon}$ is a continuous mapping, it suffices for estimations to have a limit point to have a recurrent behavior.
Lemma 1
Let $({\cal O},\mathfrak{L}_{1},\dots,\mathfrak{L}_{k},{\Upsilon})$ be a player in the environment $(\Omega^{u},{\cal F}^{u},\mathbb{P}^{u})$, and suppose there exists $(\varphi_{1}^{*},\dots,\varphi_{k}^{*})\in(\mathscr{M}_{1},\cdots,\mathscr{M}_{k})$ such that
$\mathbb{P}^{u}\Big{(}\operatorname*{liminf}_{n\to\infty}\max_{1\leq\ell\leq k}d_{\ell}(\mathfrak{L}_{\ell}({}^{n}{\cal O}),\varphi_{\ell}^{*})=0\Big{)}=1$
If ${\Upsilon}:\mathscr{M}_{1}\times\cdots\times\mathscr{M}_{k}\to({\cal D}\to{\cal P}(\mathscr{A}))$ is a continuous mapping, then ${\Upsilon}^{*}:={\Upsilon}(\varphi_{1}^{*},\dots,\varphi_{k}^{*})$ is almost surely a recurrent behavior of the player.
The main objective of this work is to show that Definition 1 is sufficiently general to encompass any player design found in the literature. A more important claim is that observed player behavior must be understood via conditions on $({\cal O},{\cal L}_{1},\dots,{\cal L}_{k},{\Upsilon})$. It is clear that some conditions on players’ observations must hold, and this is dictated by the model at hand. On the other hand, Definition 2 provides a natural condition requiring the recurrence of observed behaviors. After introducing specific estimations, we discuss how to impose conditions such as optimality, consistency, and even psychological states to better characterize both the players and the equilibrium concept.
In Section 2, we explore the setting of discrete games, introducing domains and estimations typically considered in modeling many of our games. Moreover, we introduce the concept of uncertain equilibrium and demonstrate how to refine it by imposing conditions on the estimations. Later, we provide a two-player game toy example to illustrate the dynamic nature of games even in the simplest settings. In Section 3, we explore the setting of mean-field games with constant estimations, except where the representative player estimates the population strategies. As observations can be generated by relying on symmetries, we introduce a learning algorithm with explicit examples. In Section 4, we adapt the framework from Section 2 to a single-player context and demonstrate that it can be seen as a general version of Markov Decision Processes.
@@
Literature Review. Classic non-cooperative game theory traces back to the von Neumann-Morgenstern seminal work [28] and Nash [27]. We refer to the book Maschler-Solan-Zamir [24] for a comprehensive treatment. There are many notions of equilibrium, and we highlight the correlated equilibrium introduced by Aumann [2]. The author raises common criticisms of the classic Nash equilibrium and introduces correlated equilibrium, which incorporates randomness in players’ strategies. A correlated equilibrium is a single distribution over the players’ strategies and is therefore typically motivated by a mediator who draws from this distribution. However, a single distribution cannot capture the differing knowledge each player may possess. Because of this structural connection, we will offer a comparison after introducing the uncertain equilibrium for discrete games.
@@
In our framework, players might favor having a large number of estimates, and each of them might need to be learned from observations. We cannot hope to cover every relevant learning algorithm and its literature; however, we point out the diverse work on random value functions. We also advocate that the value is fundamentally unknown to a learning player and is one of the main sources of uncertainty in planning future behavior. We refer to Thurstone [31], Luce [23], Block [6], McFadden [25], Train [32], and references therein for discussions rooted in psychology and economics. One of the oldest approaches is Thompson sampling, introduced in Thompson [30] and recently popularized by the empirical study Chapelle-Li [14]. Upper confidence bound algorithms in the context of multi-armed bandit problems, see for example Auer et al. [1], can also be viewed as a random value approach. Lastly, let us mention Bellemare-Dabney-Munos [5] who prominently promoted modeling the value function as a random variable in the context of Markov decision processes with applications to reinforcement learning.

### v4

@@ Abstract
This work introduces a unified framework for analyzing games in greater depth. In the existing literature, players’ strategies are typically assigned scalar values, and equilibrium concepts are used to identify compatible choices. However, this approach neglects the internal structure of players, thereby failing to accurately model observed behaviors.
To address this limitation, we propose an abstract definition of a player, consistent with constructions in reinforcement learning. Instead of defining games as external settings, our framework defines them in terms of the players themselves. This offers a language that enables a deeper connection between games and learning. To illustrate the need for this generality, we study a simple two-player game and show that even in basic settings, a sophisticated player may adopt dynamic strategies that cannot be captured by simpler models or compatibility analysis.
For a general definition of a player, we discuss natural conditions on its components and define competition through their behavior. In the discrete setting, we consider players whose estimates largely follow the standard framework from the literature. We explore connections to correlated equilibrium and highlight that dynamic programming naturally applies to all estimates. In the mean-field setting, we exploit symmetry to construct explicit examples of equilibria. Finally, we conclude by examining relations to reinforcement learning.
@@ 1 Introduction
The main approach of this work is to separate the concept of player from the external settings we design. This will provide a broader perspective on games, in which multiple complex decision-makers interact. In the literature, when there is a single decision-maker, designing external rewards in the environment guides which future observations the player will prefer. This, in turn, provides strong guidance for designing players that act in our interests. Still, designing players to tackle real-world problems requires significant effort beyond setting up the environment. We emphasize this perspective because once there is more than one player, the expressiveness of setting up the environment diminishes. We argue that omitting the structure of players and considering only stability conditions over the set of strategies cannot provide a sufficiently rich understanding.
Let us discuss two illustrative examples of competitive settings. The first concerns one of the simplest competitive games, rock-paper-scissors. Suppose a player competes against ten opponents in a tournament and can perfectly generate a uniform distribution over actions. Will the player choose the uniform equilibrium strategy? In that case, winning the tournament is highly unlikely. Of course, an equilibrium strategy is unexploitable, but players aiming to win the tournament would naturally avoid it. Recognizing that no one would actually use the uniform strategy, the player’s goal becomes twofold: to estimate the opponents’ behavior and simultaneously to deceive their estimates. This shows two simple but crucial points: (i) for competing players, the objective is not to be unexploitable, and (ii) the essence of the game lies not in its formal rules, but in the players themselves. Even when the setting is simple, the strategic interplay among players can exhibit profound complexity, revealing that competition lies not in the game’s structure but in its participants. The second example conveys the same ideas, but relies on the structure of the game rather than human sensory interactions. Suppose two players repeatedly change states in 0,1{0,1}, with their actions representing transition probabilities. One player aims to ”catch” the other, while the second aims to ”evade”; their rewards are defined in a zero-sum fashion. The same question arises: will they simply choose equal probabilities to remain unexploitable and earn nothing on average? In reality, if the players are truly competing, they will constantly change their strategies. For instance, the evading player might begin appearing in state $1$ more frequently to lure the opponent into following, only to deceive them for a short-term gain. Note that no equilibrium strategy yields more than zero expected gain for either side, yet that is not what the players seek. The example motivates that competition is defined not by convergence, but by the very effort to avoid it. Hence, in the next section, we will say that players are cooperating when their behaviors actually converge. As these examples show, such cooperation may occur involuntarily, when one player seeks merely to remain unexploitable.
In Section 2, we introduce the main definition of a player, and discuss some intrinsic objectives of their components. In Section 3, we explore discrete games, treating mainly the standard components as estimates of the player. After presenting specific estimates, we define uncertain equilibrium to impose conditions such as optimality and recurrence. We then refine this notion by adding further conditions such as consistency and psychological states. This allows a richer characterization of both players and the equilibrium concept. We also demonstrate connections between our optimality assumption and the concept of correlated equilibrium. Later, we present a toy two-player game to illustrate the dynamic nature of games in the simplest settings. In Section 4, we examine mean-field games with constant estimates, except that the representative player estimates the population strategies. As observations can be generated by relying on symmetries, we introduce a learning algorithm with explicit uncertain equilibrium. In Section 5, we provide further connections to core structures in reinforcement learning, present a basic learning algorithm that does not rely on standard value-based methods, and review related concepts in multi-agent reinforcement learning.
@@
Some related literature. Classic non-cooperative game theory traces back to the von Neumann-Morgenstern seminal work [89] and Nash [58]. We refer to the book Maschler-Solan-Zamir [51] for a comprehensive treatment. There are many notions of equilibrium, and we highlight the correlated equilibrium introduced by Aumann [2]. The author raises common criticisms of the classic Nash equilibrium and introduces correlated equilibrium, which incorporates randomness in players’ strategies. A correlated equilibrium is a single distribution over the players’ strategies and is therefore typically motivated by a mediator who draws from this distribution. However, a single distribution cannot capture the differing knowledge each player may possess. Because of this structural connection, we will offer a comparison after introducing the uncertain equilibrium for discrete games.
Critiques of treating equilibrium concepts as central to games appear across several literatures. Kadane and Larkey [45], for example, adopt a subjective perspective, formalizing views that had previously been discussed informally (as reflected in the Editor’s Note) and bringing them into the management science framework. In the multi-agent reinforcement learning literature, the technical note by Shoham, Powers and Grenager [72] (see also [73]) criticizes the lack of conceptual clarity arising from the unjustified use of equilibrium concepts as both learning objectives and evaluation criteria. In the behavioral game theory literature, Wright and Leyton-Brown [92] demonstrate that Nash equilibrium is a poor description of human behavior even in simple games, and provide a systematic comparison of different models of human behavior. We remark that all such models still omit a model of the human itself, mapping an external game description directly to a distribution over actions. Hartford, Wright, and Leyton-Brown [35] demonstrate that a neural network, essentially a function approximation, outperforms all such behavioral models. This strengthens the motivation for this work, as we argue that the primary modeling object should be the players themselves, and that games cannot be reduced to their external descriptions.
@@
In our framework, players may favor having a large collection of estimates, each of which must be learned from observations. We cannot hope to cover every relevant learning algorithm and its extensive literature, however, we will highlight some connections in Section 5. Here, we would like to point out the diverse work on random value functions. We also advocate that the value is fundamentally unknown to a learning player and is one of the main sources of uncertainty in planning future behavior. We refer to Thurstone [83], Luce [50], Block [9], McFadden [52], Train [84], and references therein for discussions rooted in psychology and economics. One of the oldest approaches is Thompson sampling, introduced in Thompson [82] and recently popularized by the empirical study Chapelle-Li [19]. Upper confidence bound algorithms in the context of multi-armed bandit problems, see for example Auer et al. [1], can also be viewed as a random value approach. Lastly, let us mention Bellemare-Dabney-Munos [8] who prominently promoted modeling the value function as a random variable in the context of Markov decision processes with applications to reinforcement learning.
2 Definition of Players
In this section, we provide a definition of a player and elaborate on general requirements we may impose. At the end, we also define cooperation and competition through their behavior. Let us first introduce some preliminary definitions and notation:
$\bullet$ Universe, or the environment, is an abstract probability space $(\Omega^{u},{\cal F}^{u},\mathbb{P}^{u}$);
$\bullet$ ${\cal P}(E)$ denotes the set of probability distributions on an arbitrary set $E$;
[…]
$\bullet$ $\mathbb{A}$ is the space of actions, and $\mathscr{A}$ is the set of finite sequences of $\mathbb{A}$;
$\bullet$ $\mathscr{M}_{\Upsilon}$ is the set of functions taking values in $\mathbb{A}$, called the space of behaviors;
$\bullet$ $\mathscr{M}_{\varphi}$ is the set of functions, called the space of estimates.
Similar to actions, we are keeping observables abstract. We have not yet specified the domains and ranges of the estimations. Also, $\varphi$ denotes an index for estimates. We will introduce a collection of them in the upcoming sections.
In the realm of games, a player is defined by a sequence of observations, a collection of estimates, and a sequence of actions, all of which may be highly complex. We now introduce a definition of a player:
Definition 1
We call $({\cal O},\mathfrak{L}_{\varphi},{\Upsilon})$ a player in the environment $(\Omega^{u},{\cal F}^{u},\mathbb{P}^{u})$ with observations $\cal O$}, learning algorithm $\mathfrak{L}_{\varphi}$, and with behavior $\Upsilon$}, where
${\cal O}$ $:\Omega^{u}\times\mathscr{A}\times\mathbb{N}\to\mathscr{E},$ (2.1)
$\mathfrak{L}_{\varphi}$ $:\mathscr{E}\times\mathscr{M}_{\varphi}\times\mathscr{M}_{\Upsilon}\to\mathscr{M}_{\varphi},$
${\Upsilon}$ $:\mathscr{E}\times\mathscr{M}_{\varphi}\times\mathscr{M}_{\Upsilon}\to\mathscr{M}_{\Upsilon}.$
From the perspective of mathematics, the question is what natural conditions can be imposed on this collection of functions defining a player. We begin with the consistency condition for the observations:
Definition 2
We say a player has consistent observations, if
${\cal O}(\omega^{u},a_{\cdot},n)\in\mathscr{E}\ \ \text{is a subsequence of}\ \ {\cal O}(\omega^{u},\tilde{a}_{\cdot},n+1)\in\mathscr{E}$
$\text{if}\ \ a_{\cdot}\in\mathscr{A}\ \ \text{is a subsequence of}\ \ \tilde{a}_{\cdot}\in\mathscr{A},\ \ \text{for all}\ \ n\in\mathbb{N},\omega^{u}\in\Omega^{u}$
We note that $\cal O$} sets the connection between the environment and the player, and is not available for the player to evaluate.
Next, we define a recurrence condition for a player’s behavior. To do so, we first introduce the following definition:
Definition 3
We call
$^{n}{\Upsilon}:\Omega^{u}\times\mathbb{N}\to\mathscr{M}_{\Upsilon}$
the planned behavior of the player at age $n$, where
${}^{n}{\Upsilon}$ $={\Upsilon}({}^{n}{\cal O},{}^{n}\mathfrak{L}_{\varphi},{}^{n-1}{\Upsilon})\in\mathscr{M}_{\Upsilon},$
${}^{n}\mathfrak{L}_{\varphi}$ $=\mathfrak{L}_{\varphi}({}^{n}{\cal O},{}^{n-1}\mathfrak{L}_{\varphi},{}^{n-1}{\Upsilon})\in\mathscr{M}_{\varphi},$
${}^{n}{\cal O}$ $={\cal O}(\omega^{u},{}^{n-1}I,n)\in\mathscr{E},$
${}^{n}I$ $=({}^{1}{\Upsilon}(\omega^{u},\cdot),\dots,{}^{n}{\Upsilon}(\omega^{u},\cdot))\in\mathscr{A}.$
These functions $^{n}({\Upsilon},\mathfrak{L}_{\varphi},{\cal O},I):=({}^{n}{\Upsilon},{}^{n}\mathfrak{L}_{\varphi},{}^{n}{\cal O},{}^{n}I)$ with domain $\Omega^{u}\times\mathbb{N}$ are determined in the order $^{n-1}I\to{}^{n}{\cal O}\to{}^{n}\mathfrak{L}_{\varphi}\to{}^{n}{\Upsilon}$. Behaviors in $\mathscr{M}_{\Upsilon}$ depend on $\Omega^{u}$, and evaluating at $\omega^{u}$ yields a sampled (or observed) action. That is, for $^{k}{\Upsilon}\in\mathscr{M}_{\Upsilon}$, we denote $^{k}{\Upsilon}(\omega^{u},\cdot)$ as the action taken, suppressing the rest of the unspecified domain. Let us point out that a player might have the capacity to generate randomness independently of the surrounding environment. For now, we do not explicitly track potentially independent probability spaces, such as those a player might use to sample randomness, but instead include them within the general environment. Along these lines, only the component of $\omega^{u}\in\Omega^{u}$ that is relevant to the random variable under consideration is taken into account.
We are now ready to introduce an intrinsic concept for the behavior of player. Let us equip the space $\mathscr{M}_{\Upsilon}$ with a generic metric $d$ and define:
Definition 4
We say $^{*}{\Upsilon}\in\mathscr{M}_{\Upsilon}$ is a $(r,\delta)$-recurrent behavior for a player, if
$\mathbb{P}^{u}\Big(\operatorname*{liminf}_{n\to\infty}d({}^{*}{\Upsilon},{}^{n}{\Upsilon})>r\Big)\leq\delta.$
Also, we say $^{*}{\Upsilon}$ is almost surely a recurrent behavior of the player if $r=\delta=0$.
In words, we classify behaviors that may occur infinitely often as the player ages.
Finally, let us turn to the more intricate task of imposing conditions on estimates. Motivated by the brain’s predictive nature, we introduce a notion of a player that estimates future abstract representations of observations. This lies at the core of behavior, and, roughly speaking, decision-making is about forming preferences over future observations. If we formalize observations as states, rewards, and actions, then it becomes crucial to understand the future states, rewards, and actions of other players. Then, we designate preferred future observations as those with higher total rewards. Given that observations are high-dimensional and complex, a player may need to simplify the task. For example, with visual observations, instead of predicting future pixels directly, one typically first forms useful embeddings to facilitate prediction. In case of actions, a player might estimate only intensions or goals of an oponnent. A similar reduction applies to rewards, rather than predicting future rewards directly, one may aim to learn the expected future reward. A more sophisticated agent might aim to learn a distribution of rewards.
To add structure, we introduce objects and relations formed from observations as the first layer of estimates. Let $\mathbb{N}_{\mathrm{obj}}\subset\mathbb{N}$ be an index set for different objects. For each $j\in\mathbb{N}_{\mathrm{obj}}$, let $E_{\mathrm{obj}}^{j}$ denote the space of states for object $j$. Finally, let $E_{\mathrm{rel}}$ be a space representing the set of relations. Then, let
$\mathscr{M}_{\mathrm{obj}}=\big\{\mathscr{E}\to\Pi_{j}E_{\mathrm{obj}}^{j}\big\},\ \ \mathscr{M}_{\mathrm{con}}=\big\{\Pi_{j}E_{\mathrm{obj}}^{j}\to 2^{\mathbb{N}_{\mathrm{obj}}}\big\},$
$\text{and}\quad\mathscr{M}_{\mathrm{rel}}=\big\{\Pi_{j}E_{\mathrm{obj}}^{j}\times 2^{\mathbb{N}_{\mathrm{obj}}}\to E_{\mathrm{rel}}\big\}.$
Correspondingly, we have the learning algorithms
$\mathfrak{L}_{\{{\mathrm{obj}},\mathrm{con},{\mathrm{rel}}\}}:\mathscr{E}\times\mathscr{M}_{\varphi}\times\mathscr{M}_{{\Upsilon}}\to\mathscr{M}_{\{{\mathrm{obj}},\mathrm{con},{\mathrm{rel}}\}}$
Then, $^{n}\mathfrak{L}_{\mathrm{obj}}\in\mathscr{M}_{\mathrm{obj}}$ is a mapping from observations to states of identified objects. The mapping $^{n}\mathfrak{L}_{\mathrm{con}}\in\mathscr{M}_{\mathrm{con}}$ creates connections between those objects. Finally, $^{n}\mathfrak{L}_{\mathrm{rel}}\in\mathscr{M}_{\mathrm{rel}}$ assigns relations to connections, which may take the form of discrete tags or numerical values. All of these learning algorithms may depend on past observations, estimates (denoted generically by $\varphi$), and the behavior mapping.
Now, the one-step prediction problem can be formulated as the task of modeling the next objects and relations. To represent this, let us introduce
$\mathscr{M}_{\mathrm{pre}}:=\big\{\Omega^{u}\times\Pi_{j}E_{\mathrm{obj}}^{j}\times 2^{\mathbb{N}_{\mathrm{obj}}}\times E_{\mathrm{rel}}\times\mathbb{A}\to\Pi_{j}E_{\mathrm{obj}}^{j}\times 2^{\mathbb{N}_{\mathrm{obj}}}\times E_{\mathrm{rel}}\big\}$
along with the learning algorithm $\mathfrak{L}_{\mathrm{pre}}$ and its realization $^{n}\mathfrak{L}_{\mathrm{pre}}\in\mathscr{M}_{\mathrm{pre}}$ at age $n$.
Definition 5
We say $(\varphi_{\mathrm{obj}},\varphi_{\mathrm{con}},\varphi_{\mathrm{rel}},\varphi_{\mathrm{pre}})\in(\mathscr{M}_{\mathrm{obj}},\mathscr{M}_{\mathrm{con}},\mathscr{M}_{\mathrm{rel}},\mathscr{M}_{\mathrm{pre}})$ is one-step $\varepsilon$-predictive under some metric $d$ on ${\cal P}(E_{\mathrm{obj}}\times 2^{\mathbb{N}_{\mathrm{obj}}}\times E_{\mathrm{rel}})$, if
$\hskip 0.0pt\operatorname*{liminf}_{n\to\infty}d\big(\ \mathrm{Law}({}^{n+1}E|{}^{n}{\cal O},{}^{n}I),\ \mathrm{Law}(\varphi_{\mathrm{pre}}(\omega^{u},{}^{n}E,{}^{n}{\Upsilon}(\omega^{u},\cdot))|{}^{n}{\cal O},{}^{n}I)\ \big)\leq\varepsilon,$
$\text{where}\ \ {}^{n}E=\big(\varphi_{\mathrm{obj}}({}^{n}{\cal O}),\ \varphi_{\mathrm{con}}(\varphi_{\mathrm{obj}}({}^{n}{\cal O})),\ \varphi_{\mathrm{rel}}(\varphi_{\mathrm{obj}}({}^{n}{\cal O}),\varphi_{\mathrm{con}}(\varphi_{\mathrm{obj}}({}^{n}{\cal O})))\big)$
$\mathbb{P}^{u}$-almost surely.
We note that $\varphi_{\mathrm{pre}}$ is typically a random variable whose law models the distribution of $^{n+1}E$. Let us also point out a potentially confusing notation. Given $({}^{n}{\cal O},{}^{n}I)$, the realized action $^{n}{\Upsilon}(\omega^{u},\cdot)$ is determined, whereas $\varphi_{\mathrm{pre}}(\omega^{u},\cdot)$ typically includes an independent component to model a distribution.
Let us discuss some examples to motivate the definitions above. Some related references will be provided in Section 5.
$\bullet$ We can simplify by considering the overall state of the environment as an observation, treated as a single object without further decomposition and connections. And, $\varphi_{\mathrm{rel}}$ may assign values to those states. In this case, $\varphi_{\mathrm{pre}}$ models the one-step transition probabilities of the whole state, viewed as a random variable, and the next value associated with it.
$\bullet$ Encoding the external environment is typically necessary. For humans, for example, the eyes encode visual observations from the environment. In the context of machine learning, a suitable neural network architecture carries out the encoding. Various objectives may be assigned to encoders, and $\varphi_{\mathrm{pre}}$ represents the predictive one, which might be the next embedding, a reward component of the observations, or another useful aspect of the observations for a given task.
$\bullet$ Instead of encoding the environment as a single object, one may construct many objects, each with its associated states. Then, for example, $\varphi_{\mathrm{con}}$ may form pairs, and $\varphi_{\mathrm{rel}}$ may assign attention values.
$\bullet$ Similarly, in the context of large language models (LLMs), objects can correspond to token embeddings. Here, $\varphi_{\mathrm{con}}$ forms pairs, connecting each future token to all previous ones (forming a triangular matrix), and $\varphi_{\mathrm{rel}}$ uses keys and queries to assign weights to these connections. This constitutes only a small component of such models.
$\bullet$ In the context of chess, objects can be defined as pieces together with their positions on the board. A player may then assign a large number of connections between such objects through a highly dynamic $\varphi_{\mathrm{con}}$, for example between pieces that protect each other or among larger groups representing the overall arrangement. Then, $\varphi_{\mathrm{rel}}$ may assign values to each connection, and $\varphi_{\mathrm{pre}}$ may model the behavior of the opponent.
Let us emphasize that the aim of this work is not to design new learning algorithms for estimating future observations. Our objective is to reframe common game settings from the perspective of the player and to introduce a concept of equilibrium via conditions on $({\cal O},{\cal L}_{\varphi},{\Upsilon})$. Furthermore, we emphasize that designing a player can be far more involved than setting up the game itself. We argue that understanding complex interactions between players cannot be reduced to the external setting of the game under consideration.
When there are multiple players, we keep track of them using the index $i\in\mathbb{N}_{0}:=\{1,2,\dots\}$. For any symbol $\psi$ introduced in the definitions above, we set $\vec{\psi}:=\prod_{i\in\mathbb{N}_{0}}\psi^{i}$. In this case, ${\cal E}^{i}$, the space of observables for player $i$, usually includes the actions or states of the other players. They interact and influence one another, and their collective recurrent behaviors $^{*}\vec{\Upsilon}$ form a basis for an equilibrium.
Lastly, we introduce notions of cooperation and competition defined by the players’ behavior rather than by the structure of the game itself. The goal is to separate these concepts from the external setting and use them to clarify our objectives when designing players. In essence, we ask whether the players we model are meant to exhibit stable, convergent behaviors, or to continually generate diversity through dynamic interaction.
Definition 6
We say players $(\vec{\cal O},\vec{\mathfrak{L}}_{\varphi},\vec{\Upsilon})$ in the environment $(\Omega^{u},{\cal F}^{u},\mathbb{P}^{u})$ are eventually cooperating at $^{*}\vec{\Upsilon}\in\vec{\mathscr{M}}_{\Upsilon}$, if
$\mathbb{P}^{u}\Big(\operatorname*{limsup}_{n\to\infty}\sup_{i\in\mathbb{N}_{0}}d({}^{*}{\Upsilon}^{i},{}^{n}{\Upsilon}^{i})>0\Big)=0.$
We say that players are indefinitely competing, if for any $\vec{\Upsilon}\in\vec{\mathscr{M}}_{\Upsilon}$,
$\mathbb{P}^{u}\Big(\operatorname*{limsup}_{n\to\infty}\sup_{i\in\mathbb{N}_{0}}d({\Upsilon}^{i},{}^{n}{\Upsilon}^{i})>0\Big)=1.$
As an illustration, consider a simple setting of algorithmic price competition between two firms. Although the environment is designed to be competitive, if each firm’s pricing algorithm learns that the competitor’s prices tend to move in a positively correlated manner, both algorithms can quickly converge to stable, elevated price levels. In this case, we regard the firms as cooperating, even though the underlying game remains competitive.